OpenStudy (anonymous):

A student wearing frictionless in-line skates on a horizontal surface is pushed by a friendwith a constant force of 43 N. How far must the student be pushed, starting from rest, so that her final kinetic energy is 350 J ?

OpenStudy (anonymous):

use can use work-energy theories w=theta Ke Fd=Ke2-Ke1 43*d=350-0

OpenStudy (anonymous):

okay I don't really understand how can you explain it a little more

OpenStudy (anonymous):

w=F*d newton second law F=dp/dt=d(mv)/dt we get w=(d(mv)/(dt))*d, d=v*dt w=d(m*v^2)